# EEG-DLite: Dataset Distillation for Efficient Large EEG Model Training

### v1

@@ Introduction
Large-scale EEG datasets pose new challenges to the data synthesizing approach.
Although generative approaches have shown effectiveness in computer vision tasks (zhang2024m3ddatasetcondensationminimizing; zhao2022datasetcondensationdistributionmatching), their effectiveness on physiological signals such as EEG remains largely unexplored. In our experiments, M3D, a light-weighted synthesizing framework, was selected considering the huge computational effort to generate large pretraining datasets. M3D consistently underperforms compared to all other baselines, including random sampling. As illustrated in Figure 5, the generated EEG segments exhibit unnatural characteristics, such as abrupt plateaus, flat transitions, and repetitive blocky patterns, all of which deviate from realistic brain signal dynamics. These artifacts reflect poor temporal and spectral fidelity. Furthermore, the generative process in M3D is computationally intensive, which poses scalability challenges for large-scale applications. Taken together, these findings highlight that for EEG and related physiological signals, both the synthetic quality and computational feasibility are important considerations.
Figure 6: Sample distribution per subject in SEED (seed) (top) and SEED-VII (seed-7) (bottom). Each bin represents a subject, and its height indicates the proportion of samples contributed by that subject in the original or 5% distilled dataset.
@@
BCI datasets typically involve recordings from multiple subjects, each contributing EEG segments with varying quality and characteristics. After applying diversity sampling, the percentage of samples per subject shows notable variation. As illustrated in Figure 6, some subjects contribute significantly more segments than others. This imbalance reflects inherent inter-subject variability in EEG signals, which can stem from differences in neural dynamics, recording conditions, or noise levels. Recognizing and understanding these subject-level patterns may inform several directions for future research. For instance, it may enable subject-aware pretraining strategies.
@@ Conclusion
In this study, we propose a novel data distillation framework to condense large-scale unlabeled EEG datasets through three key steps: compression, outlier removal, and diversity sampling. The framework is evaluated across four downstream tasks at varying distillation ratios, demonstrating that only 5% of the original data is sufficient to train a foundation model with comparable performance. These results highlight that data quality and diversity contribute more significantly to the generalization ability of foundation models than data quantity. We believe EEG-DLite represents a practical step toward efficient foundation model pretraining and provides insights about diverse dataset design and selection for broader downstream applications.

### v2

@@ Introduction
Large-scale EEG datasets pose new challenges to the data synthesis approach.
Although generative approaches have shown effectiveness in computer vision tasks (Zhang et al. 2024; Zhao and Bilen 2022), their effectiveness on physiological signals such as EEG remains largely unexplored. In our experiments, M3D, a light-weighted synthesizing framework, was selected considering the huge computational effort to generate large pre-training datasets. M3D consistently underperforms compared to all other baselines, including random sampling. As illustrated in Figure 5, the generated EEG segments exhibit unnatural characteristics, such as abrupt plateaus, flat transitions, and repetitive blocky patterns, all of which deviate from realistic brain signal dynamics. These artifacts reflect poor temporal and spectral fidelity. Furthermore, the generative process in M3D is computationally intensive, which poses scalability challenges for large-scale applications. Taken together, these findings highlight that for EEG and related physiological signals, both the synthetic quality and computational feasibility are important considerations.
Figure 6: Sample distribution per subject in SEED (Zheng and Lu 2015a) (top) and SEED-VII (Jiang et al. 2025) (bottom). Each bin represents a subject, and its height indicates the proportion of samples contributed by that subject in the original or 5% distilled dataset.
@@
BCI datasets typically involve recordings from multiple subjects, each contributing EEG segments with varying quality and characteristics. After applying diversity sampling, the percentage of samples per subject shows notable variation. As illustrated in Figure 6, some subjects contribute significantly more segments than others. This imbalance reflects inherent inter-subject variability in EEG signals, which can stem from differences in neural dynamics, recording conditions, or noise levels. Recognizing and understanding these subject-level patterns may inform several directions for future research. For instance, it may enable subject-aware pre-training strategies.
@@ Conclusion
In this study, we propose a novel data distillation framework to condense large-scale unlabeled EEG datasets through three key steps: compression, outlier removal, and diversity sampling. The framework is evaluated across four downstream tasks at varying distillation ratios, demonstrating that only 5% of the original data is sufficient to train a foundation model with comparable performance. These results highlight that data quality and diversity contribute more significantly to the generalization ability of foundation models than data quantity. We believe EEG-DLite represents a practical step toward efficient foundation model pre-training and provides insights about diverse dataset design and selection for broader downstream applications.